OpenStudy (anonymous):

find the speed of a planes, which accelerates from rest at 5.2m/s^2 for 6.0s. please help!

OpenStudy (anonymous):

in this question we have to find speed which can be calculated by formula:v=s/t so  now v=93.6/6=15.6 so v=15.6